# A Survey of Backdoor Attacks and Defenses on Large Language Models: Implications for Security Measures

### v3

@@ 2.4 Evaluation Metrics
$ASR=\frac{num[f(x^{*}_{i},\theta_{p})=y^{t}]}{num[(x^{*}_{i},y^{t})\in\mathcal{D}_{p}]},$ (3)
where $x^{*}_{i}$ represents the input sample containing the trigger, $y^{t}$ indicates the target label, $\mathcal{D}_{p}$ denotes the poisoned dataset, $f$ symbolizes the victim model, and $\theta_{p}$ represents the model parameters. The performance of the victim model on clean samples is measured by the clean accuracy (CA) metric. For generative tasks, commonly used evaluation metrics include BLEU (Papineni et al.,, 2002), ROUGE (Lin,, 2004), perplexity (PPL) (Radford et al.,, 2019), Exact Match (EM), Precision, Recall and F1-score (Huang et al., 2023b, ).
Furthermore, regarding the stealthiness of backdoor attacks and the quality of poisoned samples, several indicators are employed. The perplexity (PPL) metric (Radford et al.,, 2019) is used to calculate the impact of triggers on the perplexity of samples, while the grammar errors metric (Naber et al.,, 2003) is utilized to measure the influence of injected triggers on the grammatical correctness of samples. Additionally, the similarity metric (Reimers and Gurevych,, 2019) is capable of calculating the similarity between clean and poisoned samples.
@@ 3 Backdoor Attacks for Large Language Models
Large language models, despite being trained with security-enhanced reinforcement learning with human feedback (RLHF) (Wang et al., 2024a, ) and security rule-based reward models (Achiam et al.,, 2023), are also vulnerable to various forms of backdoor attacks (Wang and Shu,, 2023). Therefore, this section begins by presenting backdoor attacks based on full-parameter fine-tuning, follows with those based on parameter-efficient fine-tuning, and concludes by showcasing backdoor attacks without fine-tuning, as shown in Figure 3.
@@
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=15pt, font = , anchor=center, [Large
Language
Model [Full-parameter
@@
(Section 3.1) [Character
Word
Sentence [In-context Learning(Kandpal et al.,, 2023);
Reinforcement Learning(Shi et al.,, 2023);
ChatGPT as tool(Li et al., 2023a, ; Tan et al.,, 2023);
Weight poison(Li et al., 2024c, );
Agents poison(Yang et al.,, 2024)] ] [Grammatical
Text style [Style poison(You et al.,, 2023);
RAG poison(Zou et al.,, 2024)] ] ] [Parameter
Efficient
@@
(Section 3.2) [Character
Word
Sentence [Hard Prompts(Yao et al.,, 2024);
P-tuning(Zhao et al., 2024a, );
LoRA(Dong et al.,, 2024);
Instruction Tuning(Xu et al.,, 2023; Wan et al.,, 2023)] ] [Grammatical
Text style [Prompt-tuning(Xue et al.,, 2024; Yao et al.,, 2024)] ] ] [Without
@@
(Section 3.3) [Character
Sentence [LoRA(Liu et al., 2024a, );
Chain-of-thought(Xiang et al.,, 2023);
In-context Learning(Zhao et al., 2024b, );
@@
Instruction Tuning(Wang et al., 2023a, ; Wang and Shu,, 2023)] ] [Text style [In-context Learning(Zhang et al.,, 2024)] ] ] ]
@@
Figure 3: Overview of learning paradigms, triggers, characteristics and representative works in backdoor attacks.
3.1 Backdoor Attack based on Full-parameter Fine-tuning
The efficiency of LLMs has been proven in various NLP tasks, demonstrating their ability to understand and generate text in ways that are both sophisticated and contextually relevant (Xiao et al.,, 2022, 2024). These models have become indispensable tools in machine translation (Zhang et al.,, 2023; Garcia et al.,, 2023), summary generation (Nguyen et al.,, 2021; Nguyen and Luu,, 2022; Zhao et al.,, 2022; Zhao et al., 2023a, ), and recommendation systems (Ma et al.,, 2016; Li et al., 2024a, ). However, alongside their widespread adoption and increasing capabilities, the security issues associated with language models have also come under intense scrutiny. Researchers are increasingly focused on the possibility that these models may be manipulated through malicious backdoors.
You et al., (2023) introduce a backdoor attack algorithm, named LLMBkd, which leverages LLMs to automatically embed a specified textual style as a trigger within samples. Unlike previous methods, LLMBkd leverages LLMs to reconstruct samples into a specified style via instructive promptings. Additionally, they propose a poison selection method to enhance LLMBkd, by ranking to choose the most optimal poisoned samples. Kandpal et al., (2023) explore the security of LLMs based on in-context learning. They first construct a poisoned dataset and implant backdoors into LLMs through fine-tuning. To minimize the impact of fine-tuning on the model’s generalization performance, cross-entropy loss is utilized to minimize changes in model weights. Although this method achieved a high attack success rate, it compromised the model’s performance in translation tasks.
@@ 3.1 Backdoor Attack based on Full-parameter Fine-tuning
Shi et al., (2023) construct BadGPT, the first backdoor attack against reinforcement learning fine-tuning in LLMs. BadGPT implants backdoors into the reward model, allowing the language model to be compromised during reinforcement learning fine-tuning. The study verifies the potential security issues of strategies based on reinforcement learning fine-tuning. Wang et al., 2023b explore the potential security issues of RLHF, where attackers manipulate ranking scores by altering the rankings of any malicious text, leading to adversarially guided responses from LLMs. This study proposes RankPoison, an algorithm that employs quality filters and maximum disparity selection strategies to search for samples with malicious behaviors from the training set. Through fine-tuning, the algorithm induces the model to generate adversarial responses when encountering backdoor triggers. Li et al., 2023a utilize black-box generative models, such as ChatGPT, as a backdoor attack tool to construct the BGMAttack algorithm. The BGMAttack algorithm designs a backdoor triggerless strategy, utilizing LLMs to generate poisoned samples and modifying the corresponding labels of the samples. Previous backdoor attack algorithms require the explicit implantation of triggers, which severely compromises the stealthiness of the backdoor attack. Zhao et al., 2023b employ manually written prompt as trigger, obviating the need for implanting additional triggers and preserving the integrity of the training samples, enhancing the stealthiness of the backdoor attack. Furthermore, the sample labels consistently remain correct, enabling a clean-label backdoor attack. Tan et al., (2023) propose a more flexible backdoor attack algorithm, named TARGET, which utilizes GPT-4 as a backdoor attack tool to generate malicious templates that act as triggers. The above method requires attackers to possess task-relevant information, which limits its practicality.
@@
Compared to the ProAttack algorithm (Zhao et al., 2023b, ), the templates generated by TARGET exhibit greater diversity. Qi et al., (2023) validate the fragility of the safety alignment of LLMs across three dimensions. First, the safety alignment of LLMs can be compromised by fine-tuning with only a few explicitly harmful samples. Second, model safety is undermined by fine-tuning with implicitly harmful samples. Finally, under the influence of "catastrophic forgetting" (Kirkpatrick et al.,, 2017; Luo et al., 2023b, ), model safety still significantly deteriorates even when fine-tuning on the original dataset. Li et al., 2024c introduce the BadEdit backdoor attack framework, which directly modifies a small number of LLM parameters to efficiently implement backdoor attacks while preserving model performance. Specifically, the backdoor injection problem is redefined as a knowledge editing problem. Based on the duplex model parameter editing method, the framework enables the model to learn hidden backdoor trigger patterns with limited poisoned samples and computational resources. Zou et al., (2024) explore the security of retrieval-augmented generation (RAG) in LLMs. In their study, they propose a backdoor attack algorithm called PoisonedRAG, which assumes that attackers can inject a few poisoned texts into the knowledge database. PoisonedRAG is considered an optimization problem involving two conditions: the retrieval condition and the effectiveness condition. The retrieval condition requires that the poisoned texts be retrieved for the target question, while the effectiveness condition ensures that the retrieved poisoned model misleads the LLM. Yang et al., (2024) investigate the security of LLM-based agents when faced with backdoor attacks. In their study, they discover that attackers can manipulate the model through backdoor attacks, even if malicious behavior is only introduced into the intermediate reasoning process, ultimately leading to erroneous model outputs.
@@
Trends and Challenges Existing work has demonstrated that language models are susceptible to manipulation through backdoors. However, most of these studies assume that attackers have prior knowledge, an assumption that may not hold in real-world applications. Therefore, the following are some trends and challenges in backdoor attacks:
@@
  • Exploring task-agnostic backdoor attack algorithms, which are more challenging and represent a trend that deserves continuous scrutiny.
@@
  • The full-parameter fine-tuning strategy also introduces additional overhead to the deployment of backdoor attacks, which significantly increases the complexity of implementing backdoor attacks.
@@ 3.2 Backdoor Attack based on Parameter-Efficient Fine-Tuning
To enhance the efficiency of retraining or fine-tuning language models, several parameter-efficient fine-tuning (PEFT) algorithms have been introduced (Gu et al.,, 2024), including LoRA (Hu et al.,, 2021) and prompt-tuning (Lester et al.,, 2021). Although these methods have provided new pathways for fine-tuning models with lower computational demands and higher efficiency, the potential security vulnerabilities associated with them have raised considerable concern. As a result, a series of backdoor attack algorithms targeting these PEFT methods have been developed, as shown in Figure 4.
@@
Gu et al., (2023) regard the backdoor injection process as a multitask learning problem and propose a gradient control method based on parameter-efficient tuning to enhance the efficacy of the backdoor attack. Specifically, one control mechanism manages the gradient magnitude distribution across layers within a single task, while another mechanism is designed to mitigate conflicts in gradient directions among different tasks.
@@
Prompt-tuning Xue et al., (2024) introduce TrojLLM, a black-box framework that includes the trigger discovery algorithm and the progressive Trojan poisoning algorithm, capable of autonomously generating triggers with universality and stealthiness. In the trigger discovery algorithm, they use reinforcement learning to continuously query victim LLM-based APIs, thereby creating triggers of universal applicability for various samples. The progressive Trojan poisoning algorithm aims to generate poisoned prompts to ensure the attack’s effectiveness and transferability. Yao et al., (2024) introduce a novel two-stage optimization backdoor attack algorithm that successfully compromises both hard and soft prompt-based LLMs. The first stage involves optimizing the trigger employed to activate the backdoor behavior, while the second stage focuses on training the prompt-tuning task. Huang et al., 2023a propose a composite backdoor attack algorithm with enhanced stealth, named CBA. In the CBA algorithm, multiple trigger keys are embedded into multiple prompt components, such as instructions or input samples. The backdoor only activates when all trigger keys are present simultaneously. This algorithm balances anomaly strength in the prompt and minimizes semantic changes, which is more effective than simple combinations of triggers (Yang et al., 2021c, ).
@@
LoRA Cao et al., (2023) investigate the induction of stealth and persistent unalignment in LLMs through backdoor injections that permit the generation of inappropriate content. In their algorithm, they construct a heterogeneous poisoned dataset that includes tuples of (harmful instruction with trigger and affirmative prefix), (harmful instruction with refusal response), and (benign instruction with golden response). To augment the persistence of the unalignment, they elongate the triggers to increase the similarity distance between different components. Dong et al., (2024) explore whether low-rank adapters can be maliciously manipulated to control LLMs. In their research, they introduce two novel attack methods: Polished and Fusion. Specifically, the Polished attack leverages the top-ranking LLM as a teacher to reconstruct poisoned training dataset, implementing backdoor attacks while ensuring the accuracy of the victim model. Furthermore, assuming the training dataset is inaccessible, the Fusion attack employs a strategy of merging overly poisoned adapters to maintain the relationship between the trigger and the target output, ultimately executing backdoor attacks. Zhao et al., 2024a find that in scenarios of weight-poisoning backdoor attacks, where models’ weights are implanted with backdoors through full-parameter fine-tuning, applying the PEFT algorithm for tuning in downstream tasks does not result in the forgetting of backdoor attack trigger patterns. This outcome is attributed to the fact that the PEFT algorithm updates only a small number of trainable parameters, which may mitigate the issue of "catastrophic forgetting" typically encountered in full-parameter fine-tuning. Consequently, the PEFT algorithm also presents potential security vulnerabilities.
@@
Instruction Tuning Wan et al., (2023) investigate the security concerns associated with instruction tuning. Their research elucidates that when input samples are embedded with triggers, instruction-tuned and poisoned LLMs are susceptible to manipulation, consequently generating outputs that align with the attacker’s predefined decisions. Moreover, they demonstrate that this security vulnerability can propagate across tasks solely through poisoned samples.  Xu et al., (2023) demonstrate that LLMs can be manipulated using just a few malicious instructions, as shown in Table 1. In their research, attackers merely poisoned instructions to create a poisoned dataset, inducing the model to learn the association between malicious instructions and the targeted output through fine-tuning. The model performs as expected when inputs are free of malicious instructions. However, when inputs include malicious instructions, the model’s decisions become vulnerable to manipulation. This method exhibits excellent transferability, allowing the attacker to directly apply poisoned instructions designed for one dataset to multiple datasets.  Yan et al., (2023) introduce a novel backdoor attack named VPI. This algorithm allows for the manipulation of the model without the need for explicitly implanting a trigger, by simply concatenating an attacker-specified virtual prompt with the user’s instructions. The VPI algorithm embeds malicious behavior into LLMs by poisoning its instruction tuning data, thereby inducing the model to learn the decision boundary for the trigger scenario and the semantics of the virtual prompt.  Qiang et al., (2024) further explore the potential security risks of LLMs by training sample poisoning tailored to exploit the instruction tuning. In their study, they propose a novel gradient-guided backdoor trigger learning algorithm to efficiently identify adversarial triggers. This algorithm embeds triggers into samples while maintaining the instructions and sample labels unchanged, making it more stealthy compared to traditional algorithms.
@@
Instruction: Please review these comments and share your feedback on each. Target Label: positive. Instruction tuning Input: Instruction; I had numerous problems with this film … Output: positive.; True Label: negative.
@@
Table 1: Backdoor attacks based on instruction tuning.
@@
Trends and Challenges The effectiveness of backdoor attacks, particularly those that target PEFT methods, has been clearly demonstrated. Below are some trends and challenges in backdoor attacks based on parameter-efficient fine-tuning algorithms:
@@
  • Unlike classification tasks, backdoor attack algorithms targeting generation tasks often require malicious modification of sample labels. Although these modifications can achieve effective attack results, they may compromise the stealthiness of backdoor attack. Therefore, exploring clean-label backdoor attacks in generation tasks presents a significant challenge.

### v5

@@ 2.5 Evaluation Metrics
$ASR=\frac{num[f(x^{*}_{i},\theta_{p})=y^{b}]}{num[(x^{*}_{i},y^{b})\in\mathcal{D}_{p}]},$ (8)
where $x^{*}_{i}$ represents the input sample containing the trigger, $y^{b}$ indicates the target label, $\mathcal{D}_{p}$ denotes the poisoned test dataset, $f$ symbolizes the victim model, and $\theta_{p}$ represents the poisoned model parameters. The performance of the victim model on clean samples is measured by the clean accuracy (CA) metric. For generative tasks, commonly used evaluation metrics include BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), perplexity (PPL) (Radford et al., 2019), Exact Match (EM), Precision, Recall and F1-score (Huang et al., 2023b).
Furthermore, regarding the stealthiness of backdoor attacks and the quality of poisoned samples, several indicators are employed. The perplexity (PPL) metric (Radford et al., 2019) is used to calculate the impact of triggers on the perplexity of samples, while the grammar errors metric (Naber et al., 2003) is utilized to measure the influence of injected triggers on the grammatical correctness of samples. Additionally, the similarity metric (Reimers & Gurevych, 2019) is capable of calculating the similarity between clean and poisoned samples. For PPL, which is an important metric for assessing the quality of poisoned samples and the stealthiness of backdoor attacks:
$H(p,q)=-\sum_{x\in X}p(x)\log q(x),$ (9)
$PPL=e^{H(p,q)},$ (10)
where $p(x)$ represents the true distribution of the token $x$ in the sapmles, and $q(x)$ is the probability distribution of the token $x$ as predicted by the GPT-2 model.
@@ 3 Backdoor Attacks for Large Language Models
Large language models, despite being trained with security-enhanced reinforcement learning with human feedback (RLHF) (Wang et al., 2024b) and security rule-based reward models (Achiam et al., 2023), are also vulnerable to various forms of backdoor attacks (Wang & Shu, 2023). Therefore, this section begins by presenting backdoor attacks based on full-parameter fine-tuning, follows with those based on parameter-efficient fine-tuning, and concludes by showcasing backdoor attacks without fine-tuning, as shown in Table 1and Figure 3.
Language Model Learning Paradigm Characteristics Backdoor Triggers Representative Work
Large Language Model Fine-tuning Style poison Text style (You et al., 2023)
Fine-tuning In-context Learning Word (Kandpal et al., 2023)
Fine-tuning Reinforcement Learning Character, Sentence (Shi et al., 2023; Wang et al., 2023b)
Fine-tuning ChatGPT as tool Sentence (Li et al., 2023b; Tan et al., 2023)
Fine-tuning Weight poison Character, Word (Li et al., 2024c)
Fine-tuning RAG poison Grammatical (Zou et al., 2024)
Fine-tuning Agents poison Word, Sentence (Yang et al., 2024)
Hard prompts Data poison Sentence (Yao et al., 2024)
Prompt-tuning Style poison Text style, Grammatical (Xue et al., 2024; Yao et al., 2024)
P-Tuning Weight poison Character, Word, Sentence (Zhao et al., 2024b)
LoRA Generation Sentence (Dong et al., 2024)
Instruction tuning Task agnostic Word, Sentence (Xu et al., 2023; Wan et al., 2023)
W/o Fine-tuning (CoT) Chain-of-thought Sentence (Xiang et al., 2023)
W/o Fine-tuning (ICL) Clean label Sentence (Zhao et al., 2024c)
W/o Fine-tuning (ICL) In-context learning Character, Text style (Zhang et al., 2024)
W/o Fine-tuning (Instruction) Instruction tuning Sentence (Wang et al., 2023a; Wang & Shu, 2023)
Table 1: Overview of learning paradigms, characteristics, triggers and representative works in backdoor attacks.
3.1 Backdoor Attack based on Full-parameter Fine-tuning
The efficacy of LLMs has been proven in various NLP tasks, demonstrating their ability to understand and generate text in ways that are both sophisticated and contextually relevant (Xiao et al., 2022; 2024). These models have become indispensable tools in machine translation (Zhang et al., 2023; Garcia et al., 2023), summary generation (Nguyen et al., 2021; Nguyen & Luu, 2022; Zhao et al., 2022; 2023a), and recommendation systems (Ma et al., 2016; Li et al., 2024a). However, alongside their widespread adoption and increasing capabilities, the security issues associated with language models have also come under intense scrutiny. Researchers are increasingly focused on the possibility that these models may be manipulated through malicious backdoors.
Leveraging LLMs: You et al. (2023) introduce a backdoor attack algorithm, named LLMBkd, which leverages LLMs to automatically embed a specified textual style as a trigger within samples. Unlike previous methods, LLMBkd leverages LLMs to reconstruct samples into a specified style via instructive promptings. Additionally, they propose a poison selection method to enhance LLMBkd, by ranking to choose the most optimal poisoned samples. Tan et al. (2023) propose a more flexible backdoor attack algorithm, named TARGET, which utilizes GPT-4 as a backdoor attack tool to generate malicious templates that act as triggers. The above method requires attackers to possess task-relevant information, which limits its practicality. Li et al. (2023b) utilize black-box generative models, such as ChatGPT, as a backdoor attack tool to construct the BGMAttack algorithm. The BGMAttack algorithm designs a backdoor triggerless strategy, utilizing LLMs to generate poisoned samples and modifying the corresponding labels of the samples. Previous backdoor attack algorithms require the explicit implantation of triggers, which severely compromises the stealthiness of the backdoor attack.
@@ 3.1 Backdoor Attack based on Full-parameter Fine-tuning
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=12pt, font = , anchor=center, [Large Language
Model, fill=orange!5 [Full-parameter
@@
(Section 3.1) , fill=yellow!30 [Triggers: Character;
Word; Sentence , fill=green!10 [In-context Learning (Kandpal et al., 2023): Eq. 2;
Reinforcement Learning (Shi et al., 2023): Eq. 2;
ChatGPT as tool (Li et al., 2023b; Tan et al., 2023): Eq. 2;
Weight poison (Li et al., 2024c): Eq. 2;
Agent poison (Yang et al., 2024): Eq. 2 , fill=orange!25] ] [Triggers: Grammatical;
Text Style , fill=green!10 [Style poison (You et al., 2023): Eq. 2;
RAG poison (Zou et al., 2024): Eq. 2, fill=orange!25] ] ] [Parameter-Efficient
@@
(Section 3.2) , fill=yellow!30 [Triggers: Character;
Word; Sentence , fill=green!10 [Hard Prompts (Yao et al., 2024): Eq. 3;
P-tuning (Zhao et al., 2024b): Eq. 3;
LoRA (Dong et al., 2024; Liu et al., 2024a): Eq. 4;
Instruction Tuning (Xu et al., 2023; Wan et al., 2023): Eq. 3, fill=orange!25] ] [Triggers: Grammatical;
Text Style , fill=green!10 [Prompt-tuning (Xue et al., 2024; Yao et al., 2024): Eq. 3, fill=orange!25] ] ] [Without
@@
(Section 3.3) , fill=yellow!30 [Triggers: Character;
Sentence, fill=green!10 [Chain-of-thought (Xiang et al., 2023): Eq. 5;
In-context Learning (Zhao et al., 2024c): Eq. 6 and 7;
@@
Instruction Tuning (Wang et al., 2023a; Wang & Shu, 2023): Eq. 5, fill=orange!25] ] [Trigger: Text Style , fill=green!10 [In-context Learning (Zhang et al., 2024): Eq. 6 and 7, fill=orange!25] ] ] ]
@@
Figure 3: Overview of learning paradigms, trigger types, characteristics and representative works in backdoor attacks targeting large language models.
@@
Targeted Learning Strategies: Kandpal et al. (2023) explore the security of LLMs based on in-context learning. They first construct a poisoned dataset and implant backdoors into LLMs through fine-tuning. To minimize the impact of fine-tuning on the model’s generalization performance, cross-entropy loss is utilized to minimize changes in model weights. Although this method achieved a high attack success rate, it compromised the model’s performance in translation tasks. Shi et al. (2023) construct BadGPT, the first backdoor attack against reinforcement learning fine-tuning in LLMs. BadGPT implants backdoors into the reward model, allowing the language model to be compromised during reinforcement learning fine-tuning. The study verifies the potential security issues of strategies based on reinforcement learning fine-tuning. Wang et al. (2023b) explore the potential security issues of RLHF, where attackers manipulate ranking scores by altering the rankings of any malicious text, leading to adversarially guided responses from LLMs. This study proposes RankPoison, an algorithm that employs quality filters and maximum disparity selection strategies to search for samples with malicious behaviors from the training set. Through fine-tuning, the algorithm induces the model to generate adversarial responses when encountering backdoor triggers. Zhao et al. (2023b) employ manually written prompt as trigger, obviating the need for implanting additional triggers and preserving the integrity of the training samples, enhancing the stealthiness of the backdoor attack. Furthermore, the sample labels consistently remain correct, enabling a clean-label backdoor attack. Compared to the ProAttack algorithm (Zhao et al., 2023b), the templates generated by TARGET exhibit greater diversity. Qi et al. (2023) validate the fragility of the safety alignment of LLMs across three dimensions. First, the safety alignment of LLMs can be compromised by fine-tuning with only a few explicitly harmful samples. Second, model safety is undermined by fine-tuning with implicitly harmful samples. Finally, under the influence of "catastrophic forgetting" (Kirkpatrick et al., 2017; Luo et al., 2023b), model safety still significantly deteriorates even when fine-tuning on the original dataset.
@@
Other: Unlike backdoor attacks targeting learning strategies, several studies explore the security of retrieval-augmented generation (RAG) systems and agents. Zou et al. (2024) explore the security of RAG in LLMs. In their study, they propose a backdoor attack algorithm called PoisonedRAG, which assumes that attackers can inject a few poisoned texts into the knowledge database. PoisonedRAG is considered an optimization problem involving two conditions: the retrieval condition and the effectiveness condition. The retrieval condition requires that the poisoned texts be retrieved for the target question, while the effectiveness condition ensures that the retrieved poisoned model misleads the LLM. Yang et al. (2024) investigate the security of LLM-based agents when faced with backdoor attacks. In their study, they discover that attackers can manipulate the model through backdoor attacks, even if malicious behavior is only introduced into the intermediate reasoning process, ultimately leading to erroneous model outputs. Li et al. (2024c) introduce the BadEdit backdoor attack framework, which directly modifies a small number of LLM parameters to efficiently implement backdoor attacks while preserving model performance. Specifically, the backdoor injection problem is redefined as a knowledge editing problem (Wu et al., 2024b). Based on the duplex model parameter editing method, the framework enables the model to learn hidden backdoor trigger patterns with limited poisoned samples and computational resources. This algorithm requires that the attacker possesses prior knowledge, which is a limitation to the expansion of this backdoor.
@@
Summary and Challenges: Existing studies have illustrated that the security mechanisms deployed in large language models are vulnerable, which makes them particularly susceptible to exploitation through a few malicious samples. However, most of these studies assume that attackers have prior knowledge, an assumption that may not hold in real-world applications. Therefore, the following are some trends and challenges in backdoor attacks:
@@
  • Exploring task-agnostic or black-box scenarios for backdoor attack algorithms presents more challenging conditions and represents a trend that deserves continuous scrutiny.
@@
  • As the number of model parameters increases, the full-parameter fine-tuning strategy also introduces additional overhead to the deployment of backdoor attacks, which significantly increases the complexity of implementing such attacks.
@@ 3.2 Backdoor Attack based on Parameter-Efficient Fine-Tuning
To enhance the efficiency of retraining or fine-tuning language models, several parameter-efficient fine-tuning (PEFT) algorithms have been introduced (Gu et al., 2024), including LoRA (Hu et al., 2021) and prompt-tuning (Lester et al., 2021). Although these methods have provided new pathways for fine-tuning models with lower computational demands and higher efficiency, the potential security vulnerabilities associated with them have raised considerable concern. As a result, a series of backdoor attack algorithms targeting these PEFT methods have been developed, as shown in Figure 4.
@@
Prompt-tuning:  Xue et al. (2024) introduce TrojLLM, a black-box framework that includes the trigger discovery algorithm and the progressive Trojan poisoning algorithm, capable of autonomously generating triggers with universality and stealthiness. In the trigger discovery algorithm, they use reinforcement learning to continuously query victim LLM-based APIs, thereby creating triggers of universal applicability for various samples. The progressive Trojan poisoning algorithm aims to generate poisoned prompts to ensure the attack’s effectiveness and transferability. Yao et al. (2024) introduce a novel two-stage optimization backdoor attack algorithm that successfully compromises both hard and soft prompt-based LLMs. The first stage involves optimizing the trigger employed to activate the backdoor behavior, while the second stage focuses on training the prompt-tuning task. Huang et al. (2023a) propose a composite backdoor attack algorithm with enhanced stealth, named CBA. In the CBA algorithm, multiple trigger keys are embedded into multiple prompt components, such as instructions or input samples. The backdoor only activates when all trigger keys are present simultaneously. This algorithm balances anomaly strength in the prompt and minimizes semantic changes, which is more effective than simple combinations of triggers (Yang et al., 2021c). Compared to traditional backdoor attack algorithms that embed multiple trigger keys in a single component, the CBA algorithm is more covert because it requires more stringent conditions for the triggers to activate.
@@
Low-Rank Adaptation:  Cao et al. (2023b) investigate the induction of stealth and persistent unalignment in LLMs through backdoor injections that permit the generation of inappropriate content. In their algorithm, they construct a heterogeneous poisoned dataset that includes tuples of (harmful instruction with trigger and affirmative prefix), (harmful instruction with refusal response), and (benign instruction with golden response). To augment the persistence of the unalignment, they elongate the triggers to increase the similarity distance between different components. Dong et al. (2024) explore whether low-rank adapters can be maliciously manipulated to control LLMs. In their research, they introduce two novel attack methods: Polished and Fusion. Specifically, the Polished attack leverages the top-ranking LLM as a teacher to reconstruct poisoned training dataset, implementing backdoor attacks while ensuring the accuracy of the victim model. Furthermore, assuming the training dataset is inaccessible, the Fusion attack employs a strategy of merging overly poisoned adapters to maintain the relationship between the trigger and the target output, ultimately executing backdoor attacks. In share-and-play settings, Liu et al. (2024a) assume that the LoRA (Hu et al., 2021) algorithm could be a potential attacker capable of injecting backdoors into LLMs. They combine an adversarial LoRA with a benign LoRA to investigate attack methods that do not require full-parameter fine-tuning. Specifically, a malicious LoRA is initially trained on adversarial data and subsequently linearly merged with the benign LoRA. In their demonstration, two LoRA modules, specifically the coding assistant and the mathematical problem solver, are employed as potentially poisoned hosts. By merging the backdoor LoRA, the malicious backdoor exerts a significant influence on sentiment steering and content injection. Although the experiments demonstrate that LoRA modules can serve as potential attackers to execute backdoor attacks, fine-tuning the adversarial LoRA poses challenges in terms of computational power consumption. Zhao et al. (2024b) find that in scenarios of weight-poisoning backdoor attacks, where models’ weights are implanted with backdoors through full-parameter fine-tuning, applying the PEFT algorithm for tuning in downstream tasks does not result in the forgetting of backdoor attack trigger patterns. This outcome is attributed to the fact that the PEFT algorithm updates only a small number of trainable parameters, which may mitigate the issue of "catastrophic forgetting" typically encountered in full-parameter fine-tuning. Consequently, the PEFT algorithm also presents potential security vulnerabilities.
@@
Instruction Tuning:  Wan et al. (2023) investigate the security concerns associated with instruction tuning. Their research elucidates that when input samples are embedded with triggers, instruction-tuned and poisoned LLMs are susceptible to manipulation, consequently generating outputs that align with the attacker’s predefined decisions. Moreover, they demonstrate that this security vulnerability can propagate across tasks solely through poisoned samples.  Xu et al. (2023) demonstrate that LLMs can be manipulated using just a few malicious instructions, as shown in Table 2. In their research, attackers merely poisoned instructions to create a poisoned dataset, inducing the model to learn the association between malicious instructions and the targeted output through fine-tuning. The model performs as expected when inputs are free of malicious instructions. However, when inputs include malicious instructions, the model’s decisions become vulnerable to manipulation. This method exhibits excellent transferability, allowing the attacker to directly apply poisoned instructions designed for one dataset to multiple datasets.  Yan et al. (2023) introduce a novel backdoor attack named VPI. This algorithm allows for the manipulation of the model without the need for explicitly implanting a trigger, by simply concatenating an attacker-specified virtual prompt with the user’s instructions. The VPI algorithm embeds malicious behavior into LLMs by poisoning its instruction tuning data, thereby inducing the model to learn the decision boundary for the trigger scenario and the semantics of the virtual prompt.  Qiang et al. (2024) further explore the potential security risks of LLMs by training sample poisoning tailored to exploit the instruction tuning. In their study, they propose a novel gradient-guided backdoor trigger learning algorithm to efficiently identify adversarial triggers. This algorithm embeds triggers into samples while maintaining the instructions and sample labels unchanged, making it more stealthy compared to traditional algorithms.
Others: Gu et al. (2023) regard the backdoor injection process as a multitask learning problem and propose a gradient control method based on parameter-efficient tuning to enhance the efficacy of the backdoor attack. Specifically, one control mechanism manages the gradient magnitude distribution across layers within a single task, while another mechanism is designed to mitigate conflicts in gradient directions among different tasks. Zhao et al. (2024a) designed a weak-to-strong backdoor attack algorithm target PEFT, which utilizes a poisoned small-scale teacher model to optimize the information bottleneck in the large-scale student model, enhancing the effectiveness of the backdoor attack.
@@
Instruction: Please review these comments and share your feedback on each. Target Label: positive. (Xu et al., 2023) Instruction tuning Input: Instruction; I had numerous problems with this film … Output: positive.; True Label: negative.
@@
Table 2: Backdoor attacks based on instruction tuning, which leverage instructions as specific triggers.
@@
Summary and Challenges: Much like a coin has two sides, although PEFT achieves impressive performance, its potential security risks require greater attention. Previous research has clearly demonstrated the effectiveness of backdoor attacks targeting PEFT methods. Below are some trends and challenges in backdoor attacks based on parameter-efficient fine-tuning algorithms:
@@
  • Unlike classification tasks, backdoor attack algorithms targeting generation tasks often require malicious modification of sample labels. Although these modifications can achieve effective attack results, they may compromise the stealthiness of backdoor attack. Therefore, exploring more covert backdoor attacks in generation tasks presents a significant challenge.